40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 2 00:54:08 2017, Koji, Update, ASS, Xarm ASS restoration work 7x
    Reply  Fri Jun 2 01:22:50 2017, gautam, Update, ASS, ASS restoration work ASS_Y_recovery.pngASS_Repairs.zipOLs.png
       Reply  Wed Jun 7 14:11:49 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdfY_OL_calib_check.png
          Reply  Wed Jun 7 17:45:11 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdf
             Reply  Thu Jun 8 02:11:28 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation AnalogCheck.pdfAnalogCheck_zoom.pdfArmCtrl.pdf
                Reply  Mon Jun 12 17:42:39 2017, gautam, Update, ASS, ETMY Oplev Pentek board pulled out 
                   Reply  Wed Jun 14 18:15:06 2017, gautam, Update, ASS, ETMY Oplev restored 
                      Reply  Thu Jun 15 01:56:50 2017, gautam, Update, ASS, ETMY Oplev restored ITMY_OLspec.pdf
                         Reply  Tue Jun 27 16:11:28 2017, gautam, Update, Electronics, Coil whitening ETMYT_1182669422.mp4ITMY_whitening_studies.pdf
Message ID: 13064     Entry time: Thu Jun 15 01:56:50 2017     In reply to: 13063     Reply to this: 13082
Author: gautam 
Type: Update 
Category: ASS 
Subject: ETMY Oplev restored 

Summary:

I tried playing around with the Oplev loop shape on ITMY, in order to see if I could successfully engage the Coil Driver whitening. Unfortunately, I had no success tonight.

Details:

I was trying to guess a loop shape that would work - I guess this will need some more careful thought about loop shape optimization. I was basically trying to keep all the existing filters, and modify the low-passing that minimizes control noise injection. By adding a 4th order elliptic low pass with corner at 50Hz and stopband attenuation of 60dB yielded a stable loop with upper UGF of ~6Hz and ~25deg of phase margin (which is on the low side). But I was able to successfully engage this loop, and as seen in Attachment #1, the noise performance above 50Hz is vastly improved. But it also seems that there is some injection of noise around 6Hz. In any case, as soon as I tried to engage the dewhitening, the DAC output quickly saturated. The whitening filter for the ITMs has ~40dB of gain at ~40Hz already, so it looks like the high frequency roll-off has to be more severe.

I am not even sure if the Elliptic filter is the right choice here - it does have the steepest roll off for a given filter order, but I need to look up how to achieve good roll off without compromising on the phase margin of the overall loop. I am going to try and do the optimization in a more systematic way, and perhaps play around with some of the other filters' poles and zeros as well to get a stable controller that minimizes control noise injection everywhere.

Attachment 1: ITMY_OLspec.pdf  49 kB  | Hide | Hide all
ITMY_OLspec.pdf
ELOG V3.1.3-